OpenStudy (anonymous):

Question 4.4. Solve, using the substitution method. y + 2x = 7 14 – 4x = 2y (Points : 1) The solution is (1, 5) The solution is (21, 0) There are an infinite number of solutions. There is no solution.

OpenStudy (vishweshshrimali5):

What would be the value of y in terms of x from first equation ?

OpenStudy (vishweshshrimali5):

We have: \[\large{y + 2x = 7}\] \[\large{y = 7 - 2x}\] Any doubt ?

OpenStudy (anonymous):

no doubt @vishweshshrimali5

OpenStudy (vishweshshrimali5):

Great!! Now in second equation: \[\large{14 - 4x = 2y}\] Now, I put the value of y I obtained from eq. 1 in eq.2

OpenStudy (vishweshshrimali5):

That is: \[\large{14 - 4x = 2(7 - 2x)}\]

OpenStudy (vishweshshrimali5):

What if I simplify the RHS? That is: \[\large{2(7-2x) = 2 * 7 - 2*2x = 14 - 4x}\] Any doubt ?

OpenStudy (anonymous):

no doubt again Lol @vishweshshrimali5

OpenStudy (vishweshshrimali5):

:) So, regardless of what x is, LHS and RHS are equal. Thus: every x will satisfy the equations. So, infinite solutions